# Integration of TinyML and LargeML: A Survey of 6G and Beyond

### v1

@@ II-A1 TinyML Design
Although no standard guidelines for TinyML design currently exist, several best practices are commonly adopted to guide development.
@@
Quantization: Quantization reduces the model’s memory footprint and accelerates computation by representing parameters with lower precision numerical formats (e.g., converting 32-bit floating point values into 8-bit integers) [26]. Quantization can be implemented during training (quantization-aware training) or after training (post-training quantization). Variants include uniform and non-uniform quantization for optimizing accuracy, symmetric or asymmetric schemes to reduce quantization error, and Bayesian inference for minimizing bit encoding requirements.
@@
Pruning: Pruning reduces NN size by removing unimportant parameters, such as weights, neurons, or layers, that contribute little to model performance [27]. Pruned models typically require fine-tuning to restore accuracy. Common strategies include threshold-based pruning, in which weights below a specified magnitude are discarded. Depending on the differences in the model weight matrix, pruning is classified as unstructured (removal of individual weights) or structured (removal of neurons, filters, rows, or columns), guided by redundancy or regularization metrics.
@@
Low-rank matrix decomposition: This technique applies linear algebra methods, such as singular value decomposition or its extensions to tensors, to approximate large weight matrices with lower-rank components. The resulting factorized forms contain fewer parameters, reduce computational complexity, and simplify matrix operations [28]. Despite its benefits, this technique can be computationally intensive and more difficult to implement, as the decomposed network requires additional fine-tuning to recover accuracy loss, similar to model pruning.
@@
Weight sharing: Weight sharing reduces the number of unique parameters for storage and computation by reusing weights across different parts of the model [29]. For example, shared weights can be arranged by row or column within a weight matrix. However, designing effective sharing configurations manually is difficult and can lead to unpredictable performance, despite the inherent redundancy in DL models.
@@
Algorithm-architecture co-design: Integrating algorithm development with specialized hardware accelerators, such as TPUs or custom-designed chips (e.g., Cortex-M7, Cortex-M4, Cortex-M0+, and eDMPv1), can significantly improve the inference efficiency in TinyML models[30]. This joint optimization of algorithms and hardware also enhances computational performance and energy efficiency.
@@
Huffman coding: Huffman coding assigns variable-length codes to symbols based on their frequencies, with more frequent symbols receiving shorter codes [31]. When applied to model weights and parameters, this technique can significantly reduce overall model size through efficient encoding.
@@
Hyperparameter optimization: This refers to the process of selecting the optimal configuration of model parameters, such as learning rate, batch size, activation functions, number of neurons, and network depth [32]. Common optimization methods include grid search, random search, Bayesian optimization, Hyperband (which dynamically allocates resources to promising configurations while terminating underperforming configurations early), genetic algorithms, tree-structured Parzen estimators, and hyperdimensional computing. These techniques aim to improve model performance by minimizing memory and computational overhead while ensuring user-defined accuracy.
@@
Knowledge distillation: Knowledge distillation involves transferring learned representations from a larger model (the teacher) to a smaller model (the student) to produce a compact model with comparable performance. Three common types of distillation techniques are (1) response-based, where the student is trained to mimic the teacher’s final output layer, (2) feature-based, where the student replicates intermediate feature representations from the teacher, and (3) relation-based, where the student learns the structural relationships between different components of the teacher model [33].
@@ II-A2 TinyML Operation
Firmware updates: This involves replacing the entire software stack on a device, including the TinyML model and underlying system components. Although comprehensive, this process is resource-intensive and time-consuming. Firmware updates are typically delivered over-the-air (OTA) using lightweight machine-to-machine (M2M) protocols, which support both push and pull modes for firmware transfer [35]. Tools such as RIOT-ML facilitate secure OTA updates, enabling remote, unattended deployment.
Model operations: LargeML deployment involves multiple strategies to optimize model updates and execution on resource-constrained devices [34]. Full model updates replace the entire ML model via OTA delivery, ideal for significant architecture or parameter changes, ensuring devices run the latest version. Partial model updates target specific layers or parameters, minimizing network traffic and power use, making them suitable for fine-tuning on constrained devices [36]. Model transpilers convert models between frameworks, like TensorFlow to TensorFlow Lite Micro, enhancing compatibility and resource efficiency across diverse hardware [37]. Model compilers like uTensor and TinyTS translate models into optimized binary code for microcontrollers, using quantization and pruning to reduce size and boost inference efficiency, fitting within tight memory and computational limits.
Communication protocols: TinyML systems leverage specialized communication protocols to optimize data transfer under power, bandwidth, and range constraints in IoT ecosystems [15]. For intra-device communication, protocols like inter-integrated circuit and serial peripheral interface enable efficient data exchange between microcontrollers and peripherals during local collection. Device-to-cloud communication often employs message queuing telemetry transport, a lightweight protocol ideal for low-bandwidth, high-latency environments. Short-range wireless needs are met by low-power wide-area network (LoWPAN) protocols such as Bluetooth, Zigbee, or 6LoWPAN for wearables and smart home sensors, while Wi-Fi supports higher data volumes in smart cameras or environmental monitors. For long-range, low-power communication, long-range wide area network excels in remote applications like agriculture and smart cities, offering kilometer-scale coverage with minimal energy use. Sigfox, designed for ultra-narrowband IoT and M2M tasks, transmits small, infrequent data packets on unlicensed bands, ensuring low-cost, secure deployments. Similarly, Narrowband IoT, a 3GPP standard, supports low-data-rate, energy-efficient communication in licensed/unlicensed spectra, enabling scalable, reliable connectivity for massive IoT deployments in challenging environments like underground or indoor settings.
@@ II-A3 Performance Evaluation
Deployment robustness: Deployment robustness ensures TinyML model reliability across environmental conditions. Two evaluation approaches are used [34]: per-model evaluation assesses performance in controlled pre-deployment tests, while per-operator evaluation monitors adaptability in real-world settings. Robustness testing involves diverse datasets, hardware, and environments to ensure consistent performance.
II-B LargeML
Integrating LargeML into wireless networks is a key trend for enabling autonomous operations, using cognitive modules for resource allocation, channel estimation, and mobility management [5]. The LargeML life cycle consists of three stages:
  • Stage 1: Collect and pre-process datasets from multiple domains to construct a high-quality training corpus.
  • Stage 2: Design model architectures by selecting appropriate layer types and arrangements, configuring parameters, and determining suitable training algorithms.
  • Stage 3: Train the model in two phases: (1) Pre-train on broad datasets to build generalized representations; (2) Fine-tune on task-specific data for applications such as question answering, summarization, and code generation.
  • Stage 4: Evaluate performance using established benchmarks. Refine model architecture, hyperparameters, and datasets based on results.
  • Stage 5: Apply model compression techniques to reduce the memory footprint and inference latency without sacrificing performance.
  • Stage 6: Test and optimize the model in controlled environments to maximize inference efficiency without degrading output quality.
  • Stage 7: Deploy in production settings. Continuously monitor, fine-tune, and adapt the model to accommodate evolving data patterns and user requirements.
  • Stage 8: Mitigate risks and enhance transparency by filtering outputs through techniques such as adversarial training, fairness constraints, and curated datasets. These steps address bias, privacy, security, and misuse concerns.
@@ II-B1 LargeML Design
The design process for LargeML covers six key areas: data processing, learning architecture, network configuration, pre-training tasks, fine-tuning strategies, and prompt engineering. Each area is outlined below.
@@
Training LargeML relies on extensive IoT filme data, where data quality significantly impacts prediction and decision accuracy. Effective data processing involves key actions to ensure high-quality outcomes.
Collection and labeling: Data is gathered from heterogeneous sources (i.e., real-time sensors, network logs, user feedback, and environmental inputs). These data points are annotated with application-relevant features (e.g., signal strength, noise levels, and user mobility patterns [40]). Consistent and accurate labeling is critical to prevent bias in state or event label representations [41]. Well-structured datasets enable the model to generalize across a range of operational scenarios.
Cleaning, transformation, and augmentation: This phase prepares the raw data for modeling. Cleaning involves removing duplicates, correcting errors, handling missing values, and filtering noise. Transformation includes normalization, encoding categorical variables (e.g., one-hot or label encoding), and feature engineering. These transformations convert raw data into model-ready formats for analysis, such as token-based input (for textual and code sequences), tree or graph-based structures (to capture syntactic or hierarchical relationships), pixel-based representations (for visual encoding of data), and hybrid formats that integrate multiple representations [17]. Augmentation techniques, including rotation, flipping, or noise injection, generate additional samples to enhance model robustness. The overall process ensures a clean, consistent, and well-structured dataset for downstream tasks.
Featuring and splitting: This phase involves feature selection, feature extraction, and data partitioning [42]. Feature selection identifies the variables with the greatest impact on model performance, a process often referred to as sensitivity analysis. Feature extraction derives new variables, for example the rate of change in wireless signal strength, to reveal deeper patterns within the data. After the features are defined, the dataset is partitioned into training, validation, and testing sets to evaluate the model’s performance and prevent bias. Cross-validation, involving training the model on different subsets of the data, is used to assess robustness and generalizability.
@@
Mixture of experts (MoE): This technique extends all three architectural types by scaling a subset of NN weights, enabling sparse activation. Models such as Switch Transformer [46] and GLaM [47] demonstrate MoE’s ability to increase model parameter counts without proportional increases in computational cost. Performance improves with more experts or larger total parameter size. However, training can become unstable due to complex routing operations. To address this, techniques such as incorporating high-precision tensors in the routing module, narrowing initialization ranges, or introducing resource-conditional computation alongside traditional data-conditional computation have shown promise [48].
Emergence: Conventional transformer architectures may struggle with quadratic computational complexity with respect to sequence length, resulting in high processing costs for long inputs. This limitation has prompted research into parameterized state space architectures [49], which integrate features of RNNs and CNNs. These models generate outputs recursively in the same manner as RNNs, referencing only the previous state during decoding while avoiding the need to revisit all previous states, which is typical in conventional transformers. These models also encode entire sequences in parallel using convolution operations, exploiting GPU efficiency through techniques such as parallel scan, fast Fourier transform, and chunkwise recurrent processing.
@@
Network configuration directly impacts model training and performance. Selecting the appropriate operational strategy for the model is a critical early step. The following approaches highlight important considerations for efficient network design [50].
Parallel training: This technique accelerates training through parallelism strategies. Data parallelism synchronizes GPUs using a parameter server/all-reduce gradients. Pipeline parallelism divides the model across devices to overlap the execution of data processing. Model parallelism distributes layers or neurons across devices, with each device processing parts of the model and exchanging intermediate outputs.
Mixed precision: During model training, parameter values typically remain within the range of 16-bit floating-point numbers. However, the product of gradients and the learning rate during updates can exceed this range, leading to underflow and ineffective parameter updates. To prevent this problem, optimizers store an auxiliary set of parameters in single-precision (32-bit). While training uses half-precision (16-bit) for parameters and gradients, the optimizer computes and stores updates in 32-bit before conversion into 16-bit.
Offloading: Optimizer parameters often consume more memory than model parameters, requiring offloading from GPU to CPU. However, offloading to a single CPU reduces the training throughput. This bottleneck is overcome by distributing the load across multiple CPUs linked to the GPU.
Overlapping: Memory operations are generally asynchronous. To improve computational efficiency during the model’s forward propagation, a two-phase operation is applied. Memory requests (parameter retrieval) are issued first, followed by computations (parameter processing), allowing the memory operation to complete (new parameter collection).
Checkpointing: In transformer models, backward propagation uses a checkpoint mechanism to optimize memory by saving only selected inputs (checkpoints) at each major layer (attention and feed-forward) during forward propagation, rather than storing all intermediate results. During backpropagation, gradients are computed by re-executing the forward pass for each layer using these checkpoints, after which both checkpoints and recomputed results are discarded, balancing memory efficiency with computational cost.
@@
Pre-training establishes the groundwork for the model by equipping it with essential skills and knowledge that require minimal task-specific training, involving extensive data collection and preparation, self-supervised learning to acquire foundational knowledge, and fine-tuning for specific tasks. Before deployment, LargeML undergoes pre-training on extensive corpora to learn universal representations that encode both syntactic and semantic information. The models are then fine-tuned for downstream tasks, facilitating knowledge transfer to a broad range of applications. Below are common pre-training workflows [51].
Single modal: This method trains models on a single data modality (e.g., text, images, audio) to build domain-specific representations. In NLP, for example, pre-training on large text corpora enables models to capture linguistic structures and semantics, excelling in tasks such as sentiment analysis, speech signal classification, and machine translation.
Multiple modal: This method combines different data types (e.g., text, images) to build richer representations. For example, training on both image captions and corresponding images enables models to link visual and textual information, improving performance in tasks such as image captioning, visual question answering, and cross-modal retrieval.
Token prediction: Token prediction tasks train models to predict the next token in a sequence, strengthening contextual interpretation and language processing. Masked language modeling (e.g., BERT) hides random tokens for prediction, while autoregressive modeling (e.g., GPT) predicts the next token based on prior content, supporting coherent and contextually accurate outputs.
@@
After pre-training, a model is fine-tuned on a smaller, task-specific dataset to adjusts its parameters for the target application. This process builds on the knowledge acquired during pre-training. In addition to the TinyML strategies discussed in Section II-A1, this section reviews the fine-tuning techniques relevant to LargeML systems [52].
@@
Low-rank adaptation (LoRA): LoRA enhances transformer models by introducing trainable low-rank matrices to selected layers, keeping original weights frozen [54]. Only these lightweight matrices are trained to approximate weight updates, enabling efficient fine-tuning with minimal memory and computational costs.
Knowledge preservation: This technique adapts pre-trained models to new tasks while retaining their valuable pre-existing knowledge. It addresses catastrophic forgetting, using techniques such as causal graph framing to preserve important causal effects from pre-trained data, and PEFT techniques such as LoRA and Half Fine-Tuning. This balance ensures that the model remains effective across diverse tasks without overfitting to new data.
@@
After fine-tuning, prompt engineering refines how the model interprets and responds to inputs. Prompting techniques vary depending on task requirements and application contexts. Common types include multi-turn instructions, multiple modeling, task planning, tool augmented alignment, and retrieval augmentation. The most widely used prompting techniques in wireless networks are described below [5].
Few-shot prompting: This method enables the model to generalize from a small number of examples. It is particularly useful when large annotated datasets are unavailable, allowing the model to perform new tasks with minimal additional data.
Zero-shot prompting: This method exploits the model’s pre-trained knowledge to perform tasks without prior examples or specific training. The model generates outputs based solely on instructions, drawing from its existing knowledge. In contrast to few-shot prompting, zero-shot prompting provides no examples to guide the model’s output.
In-context learning: This method refers to a model’s ability to infer patterns and peform tasks based solely on prompts, without modifying its internal parameters. The user defines the task and provides examples within the prompt. The model uses attention mechanisms to analyze these examples and deduce the task structure. The model then applies the inferred pattern to generate outputs for new inputs. For example, LargeML can use in-context learning for network intrusion detection, adaptive response to threats, or for resource management, optimizing resource allocation in real time.
Chain of thought: This approach improves reasoning by breaking complex tasks into intermediate steps. It guides the model to generate rationales for each step, processing information logically and maintaining contextual coherence. Consequently, each decision is based on established information. A practical application is fault diagnosis and troubleshooting, where LargeML analyzes network issues step-by-step, accelerating resolution and improving system resilience.
Prompt-based learning: This method uses the model’s pre-trained knowledge to generate accurate, contextually appropriate responses without additional training or parameter adjustments. Prompts focus the model’s attention and shape its output for specific tasks. For example, LargeML can analyze historical data to predict maintenance needs and preempt outages by identifying failure patterns.

### v2

@@ II-A1 TinyML Design
While no standard design guidelines exist, several best practices are applied to guide development.
@@
Quantization: Reduces the memory footprint and computation by representing parameters in lower precision (e.g., 8-bit integers versus 32-bit floats) [144]. It can be applied during training (quantization-aware training) or after training (post-training quantization). Variants include uniform and non-uniform quantization for optimizing accuracy, symmetric or asymmetric schemes to reduce quantization error, and Bayesian inference to minimize bit encoding requirements.
@@
Pruning: Removes redundant weights, neurons, or layers to shrink NN size [212]. Pruned models typically require fine-tuning to restore accuracy. Common strategies include threshold-based pruning, in which weights below a specified magnitude are discarded. From differences in the model weight matrix, pruning is classified as unstructured (removal of individual weights) or structured (removal of neurons, filters, rows, or columns), guided by redundancy or regularization metrics.
@@
Low-rank matrix decomposition: Approximates large weight matrices via techniques like singular value decomposition, with lower-rank components. The resulting factorized forms contain fewer parameters, reduce computational complexity, and simplify matrix operations [171]. Albeit this technique can be computationally intensive and more difficult to implement, as the decomposed network requires additional fine-tuning to recover accuracy loss, similar to model pruning.
@@
Weight sharing: Reduces parameters across model components to cut storage and computation [126]. For example, shared weights can be arranged by row or column within a weight matrix. However, designing effective sharing configurations is difficult and can lead to unpredictable performance.
@@
Algorithm-architecture co-design: Jointly optimizes algorithms with specialized hardware (e.g., TPUs, Cortex-M7, Cortex-M4, Cortex-M0+, and eDMPv1), for inference efficiency [87]. This joint optimization also enhances computational performance and energy efficiency.
@@
Huffman coding: Compresses model parameters using variable-length encoding based on symbol frequency [116]. When applied to model weights and parameters, this technique can reduce overall model size through efficient encoding.
@@
Hyperparameter optimization: Tunes model parameters, such as learning rate, batch size, activation functions, number of neurons, and network depth, typically via grid search, random search, Bayesian optimization, Hyperband (which dynamically allocates resources to promising configurations while terminating underperforming configurations early), genetic algorithms, tree-structured Parzen estimators, and hyperdimensional computing [47]. These techniques aim to improve model performance by minimizing memory and computational overhead while ensuring user-defined accuracy.
@@
Knowledge distillation: Transfers knowledge from a larger “teacher” model to a smaller “student” via (i) response-based, where students mimic the teacher’s final output; (ii) feature-based, where students replicate intermediate feature representations; and (iii) relation-based, where students learn the structural relationships within the teacher model [194].
@@ II-A2 TinyML Operation
Firmware updates: Replace the entire software stack, including the TinyML model and system components. While comprehensive, this process is resource- and time-intensive. Over-the-air (OTA) delivery via lightweight M2M protocols (push/pull) is common [155]. Tools such as RIOT-ML facilitate secure OTA updates, enabling remote, unattended deployment.
Model operations: Several strategies optimize TinyML deployment [66]. Full model updates replace the entire ML model via OTA delivery, ideal for major changes, ensuring devices run the latest version. Partial model updates modify only specific layers or parameters, reducing network traffic and power use, suitable for fine-tuning on constrained devices [125]. Model transpilers (e.g., TensorFlow $\to$ TensorFlow Lite Micro), improve cross-hardware compatibility and resource efficiency [37]. Model compilers (e.g., uTensor, TinyTS) generate optimized binaries using quantization and pruning, reducing size and improving inference efficiency, fitting within tight memory and computational limits.
Communication protocols: TinyML relies on specialized protocols to optimize data transfer under hardware constraints in IoT ecosystems [23]. For intra-device links, inter-integrated circuit and serial peripheral interface protocols support efficient data exchange between MCU-peripheral exchange. For device-cloud links, message queuing telemetry transport offers lightweight transfer in low-bandwidth, high-latency environments. Short-range links are met by low-power wide-area network (LoWPAN) protocols (Bluetooth, Zigbee, and 6LoWPAN) for wearable and smart sensors, while Wi-Fi supports higher data rates. For long-range, low-power links, long-range wide area network excels in remote applications like agriculture and smart cities, offering energy-efficient, kilometer-scale coverage. Sigfox, designed for ultra-narrowband IoT and M2M tasks, transmits small, infrequent data packets on unlicensed bands, ensuring low-cost, secure deployments. Similarly, Narrowband IoT (3GPP) supports low-data-rate, energy-efficient communication in licensed/unlicensed spectra, enabling scalable, reliable connectivity for massive IoT deployments in challenging environments like underground or indoor settings.
@@ II-A3 Performance Evaluation
Deployment Robustness: Deployment durability ensures the reliability of the TinyML model under a variety of environmental conditions. Two evaluation methods are used [66]: per-model evaluation assesses performance in controlled pre-deployment tests, while per-operator evaluation monitors adaptability in real-world environments. Durability testing encompasses diverse datasets, hardware, and environments to ensure consistent performance.
II-B Overview of LargeML
LargeML is increasingly integrated into wireless networks to enable autonomous operations via cognitive modules for resource allocation, channel estimation, and mobility management [193]. Its life cycle consists of eight stages: (i) Stage 1: Data collection and pre-processing from multiple domains for high-quality training corpus; (ii) Stage 2: Model architecture design regarding layer types, parameters, and training algorithms; (iii) Stage 3: Training through pre-training on broad data and fine-tuning for task-specific applications; (iv) Stage 4: Evaluation with benchmarks, refining architecture, hyperparameters, and datasets; (v) Stage 5: Compression to reduce memory and latency while retaining performance; (vi) Stage 6: Testing and optimization under controlled conditions for inference efficiency; (vii) Stage 7: Deployment with continuous monitoring, fine-tuning, and adaptation; and (viii) Stage 8: Risk mitigation using adversarial training, fairness constraints, and curated datasets, addressing bias, privacy, security, and misuse concerns.
@@ II-B1 LargeML Design
The design process for LargeML covers six key areas, outlined below.
@@
Training LargeML models depends on diverse IoT data, as quality impacts prediction accuracy. The process involves three key steps.
Collection and labeling: Data is gathered from heterogeneous sources (real-time sensors, logs, user feedback) and annotated with relevant features (e.g., signal strength, noise, mobility patterns [86]). Accurate labeling prevents bias and ensures models generalize across operational scenarios [63].
Cleaning, transformation, and augmentation: Raw data is cleaned by removing duplicates, correcting errors, and handling missing values. It is then transformed through normalization, encoding, and feature engineering, and represented in formats suitable for models, such as token-based input for text, tree structures for hierarchical relationships, and pixel-based representations for visual data [24]. Augmentation methods, like rotation, flipping, and noise injection, enhance sample diversity and robustness, resulting in a clean and well-structured dataset for downstream tasks.
Featuring and splitting: It involves feature selection to identify high-impact variables, feature extraction to derive new indicators (e.g., signal variation rates) and capture deeper patterns, as well as data partitioning for training, validation, and test sets to ensure robustness and generalizability [119].
@@
Mixture of experts (MoE): MoE activates only subsets of model weights, improving efficiency as in Switch Transformer [45] and GLaM [41]. Performance scales with more experts or larger parameter size, but unstable training can occur due to complex routing operations. This issue can be mitigated by precision tensors, initialization adjustments, and hybrid resource-conditional computation [166].
Emergence architectures: Standard transformers face quadratic complexity with sequence length. Motivated by this, parameterized state-space models [53] integrate RNN- and CNN-like properties: they process outputs recursively (like RNNs) while encoding sequences in parallel (like CNNs). Techniques such as fast Fourier transform and chunk-wise recurrence can further improve efficiency.
@@
Network setup strongly affects training performance. Common strategies address parallelism, precision, memory, and computation [107].
Parallel training: Data parallelism synchronizes GPUs using a parameter server/all-reduce gradients; pipeline parallelism splits models across devices; and model parallelism distributes layers/neurons, exchanging intermediate results.
Mixed precision and offloading: Mixed precision stores gradients in single-precision (32-bit) while training in half-precision (16-bit) to prevent underflow and ineffective parameter updates. Offloading optimizer states from the GPU to multiple CPUs alleviates memory bottlenecks.
Overlapping and check-pointing: Overlapping fetches parameters while computing on previously loaded ones, and checkpointing stores only minimal activations and recomputes the rest to trade memory for compute.
@@
Pre-training equips models with general knowledge that minimizes the need for task-specific training. It involves large-scale data collection, self-supervised learning to capture foundational representations, and fine-tuning for target tasks. LargeML models are pre-trained on extensive corpora to learn universal syntactic and semantic features, which are later adapted for downstream applications. Common pre-training workflows are outlined below [186].
Single-modal: Trains on a single modality (e.g., text, images, audio) to capture domain-specific patterns. For instance, pre-training on large text corpora enables strong performance in sentiment analysis, speech classification, and translation.
Multi-modal: Combines different modalities to learn richer representations. For example, training image captions with visuals improves tasks such as captioning, visual question answering, and cross-modal retrieval.
Token prediction: Trains models to predict missing or next tokens, enhancing contextual understanding. Masked language modeling (e.g., BERT) predicts hidden random tokens, while auto-regressive modeling (e.g., GPT) generates tokens sequentially for coherent outputs.
@@
After pre-training, a model is fine-tuned on a smaller, task-specific dataset to adjust its parameters for the target application. This process builds on the knowledge acquired during pre-training. In addition to the TinyML strategies discussed in Section II-A1, this section reviews the fine-tuning techniques relevant to LargeML systems [60].
@@
Low-rank adaptation (LoRA): LoRA enhances transformer models by introducing trainable low-rank matrices to selected layers, keeping original weights frozen [62]. Only lightweight matrices are trained to approximate weight updates, thereby reducing memory and computational costs.
Knowledge preservation: This technique adapts a pre‑trained model to new tasks while preserving its valuable pre-existing knowledge, mitigating catastrophic forgetting through causal‑aware methods and PEFT techniques such as LoRA and Half Fine‑Tuning. The result is a model that stays effective across tasks without overfitting to new data.
@@
After fine-tuning, prompt engineering refines how the model interprets and responds to inputs according to task requirements and application contexts. Common types include multi-turn instructions, multiple modeling, task planning, tool augmented alignment, and retrieval augmentation. The most widely used prompting techniques in wireless networks are described below [193].
Few-shot prompting: Enables the model to generalize from a small number of examples. It is particularly useful when large annotated datasets are unavailable, allowing the model to perform new tasks with minimal additional data.
Zero-shot prompting: Exploits the model’s pre-trained knowledge to perform tasks without prior examples or specific training. The model generates outputs based solely on instructions, drawing from its existing knowledge. In contrast to few-shot prompting, zero-shot prompting provides no examples to guide the model’s output.
In-context learning: Enables models to perform tasks from prompts alone, without parameter updates. By analyzing examples in the prompt, the model infers task structure and applies it to new inputs, e.g., detecting network intrusions, adapting to threats, or managing resources in real time.
Chain-of-thought: Enhances reasoning by breaking tasks into intermediate steps, ensuring logical and coherent decision-making. Useful in applications where stepwise analysis accelerates repairing and improves resilience, e.g., fault diagnosis.
Prompt-based learning: Leverages pre-trained knowledge to generate accurate, contextually relevant responses without further training or parameter adjustments. Well-designed prompts guide task execution, e.g., predicting maintenance needs and preventing outages by identifying failure patterns.